# Nonparametric Uniform Inference in

### v1

@@ 6.3 Bootstrap Inference for the Optimal Policy Value
6.4 An Example of $\widehat{g}$ and Sufficient Conditions
Thus far in Section 6, we have allowed for an arbitrary preliminary estimator $\widehat{g}$ of the surrogate policy $g_{\ast}$, provided that it satisfies the high-level condition in Assumption 6.1. The current subsection shows that the cross-fitted sieve estimator $\widehat{g}$ proposed in Section 5.1 can be used for all purposes of estimation and inference for the optimal policy value developed in Section 6, since it satisfies the high-level condition in Assumption 6.1, as formally stated below.
Proposition 1:
If the conditions of Theorem 5.1 hold and $k^{4+\varepsilon}/n\lesssim 1$ for some constant $\varepsilon>0$, then the cross-fitting sieve estimator $\widehat{g}$ of (5.1) satisfies Assumption 6.1.
This proposition formally endorses the validity of the cross-fitted sieve estimator $\widehat{g}$ from Section 5.1. In addition, other commonly used estimators, such as the full-sample sieve estimator or the $\ell_{1}$-penalized estimator, can be shown to satisfy the same condition by analogous arguments. For succinctness, the details of these alternative estimation strategies are omitted here. Moreover, parametric estimators can readily meet the requirement by treating the sieve space $\mathcal{G}_{k}$ as the parametric class $\mathcal{G}$, in which case the sieve approximation error entirely vanishes.
@@ 7 Simulation Studies
The surrogate policy function is estimated using a sieve approximation based on a tensor-product Legendre polynomial basis. Specifically, for $k\in\{2,3,4\}$, we employ the tensor-product basis constructed from the orthonormal Legendre polynomials $\{p_{j}(x)\}_{j=0}^{\infty}$ of $k$-dimensions defined on the support $\mathcal{X}$. The sieve basis is given by $p_{1:k}(x_{1},x_{2})=\big\{\,p_{j}(x_{1})\,p_{l}(x_{2}):j,l\in[k]\,\big\},$ and the policy function is approximated as $g_{k}(x_{1},x_{2})=\sum_{j=1}^{k}\sum_{l=1}^{k}\beta_{jl}\,p_{j}(x_{1})p_{l}(x_{2}),$ or, equivalently, $g_{k}(x_{1},x_{2})=p_{1:k}(x_{1},x_{2})^{\prime}\beta.$
@@ 7.2 Gaussian Limit and Inference for the Optimal Welfare
Each plotted density from (a) to (d) closely resembles the standard normal distribution. In particular, the result in panel (a) for the plug-in estimated optimal welfare corroborates the theoretical property of Gaussian limit behavior established in Theorem 6.2. Panel (a) is comparable to panel (b) for the oracle counterpart, and this observation is consistent with the asymptotic negligibility of the preliminary surrogate policy estimation emphasized in Section 6.2. Although we report one set of results based on the logistic loss, these findings hold robustly across different surrogate loss functions.
@@ 9 Conclusion
In this paper, we have developed a toolkit for nonparametric uniform inference in cost-sensitive binary classification, a broad framework that encompasses maximum score estimation, utility-maximizing choice prediction, policy learning, and related problems. These settings are well known for slow convergence and nonstandard limiting behavior, and nonparametric formulations of classification policy functions introduce the additional challenge of failures of identification.
To address these difficulties, we proposed a strictly convex surrogate loss that point-identifies a canonical nonparametric classification policy function. Estimating this surrogate policy function provides a unified path to inference for both the optimal classification rule and its associated policy value. In particular, this strategy delivers a Gaussian approximation for the optimal classification policy and a Gaussian limit for the optimal policy value, enabling standard inferential procedures and substantially simplifying empirical implementation despite the intrinsic difficulty of the underlying problem.
Simulation evidence corroborates the theoretical results, showing that the proposed methods perform well in finite samples. Finally, our application to the National JTPA Study demonstrates how the procedure can be used to recover an optimal treatment assignment rule and to evaluate its welfare implications.
@@ A.1 Proof for Section 4: Surrogate Identification
$C_{1}(x)=\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}=x]\quad\text{and}\quad C_{0}(x)=\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}=x].$
@@
Note that $g_{\lambda}\in\mathcal{G}$ because $\mathcal{G}$ is convex. Since $\phi$ is strictly convex, $\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}]>0$ a.s., and $\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}]>0$ a.s.,
$\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}]\cdot\phi(-g_{\lambda}(X_{i}))+\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}]\cdot\phi(g_{\lambda}(X_{i}))$
$\leq$ $\lambda[\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}]\cdot\phi(-g_{1}(X_{i}))+\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}]\cdot\phi(g_{1}(X_{i}))]$
$+(1-\lambda)[\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}]\cdot\phi(-g_{2}(X_{i}))+\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}]\cdot\phi(g_{2}(X_{i}))],$
@@ A.3 An Auxiliary Result Related to Section 6
$C_{1}(x)=\mathbb{E}[\psi^{+}(Z_{i},\eta)|X_{i}=x]\quad\text{and}\quad C_{0}(x)=\mathbb{E}[\psi^{-}(Z_{i},\eta)|X_{i}=x].$
@@
Suppose Assumptions 5.3, 6.2, 6.3.(ii), and conditions of Theorem 4.1(ii) hold. Then, there exists small $\delta$ such that for any $\varepsilon<\delta$, $\mathbf{1}\{|g_{\ast}(x)|\leq\varepsilon\}\leq\mathbf{1}\{|C_{1}(x)-C_{0}(x)|\leq K_{1}\varepsilon\}$ for a positive $K_{1}$ that does not depends on $x$ and $\varepsilon$.
@@
In addition to the conditions for Theorem 4.1(ii), suppose that that Assumptions 6.1 to 6.3 hold. Then, we have
@@
We now analyze the second term on the right-hand side. Note $|V(\widehat{g})-V(g_{k})|\leq|V(\widehat{g})-V(g_{\ast})|+|V(g_{\ast})-V(g_{k})|$. Let $U_{i}:=\mathbb{E}[\psi_{1}(Z_{i},\eta)|X_{i}]-\mathbb{E}[\psi_{0}(Z_{i},\eta)|X_{i}]$. Then, with some positive constant $K_{1}$,
@@
where the first inequality holds because $\widehat{g}(X_{i})$ and $g_{\ast}(X_{i})$ disagree implies that $|g_{\ast}(X_{i})|\leq|\widehat{g}(X_{i})-g_{*}(X_{i})|$, and when they disagree, the absolute difference $|V(\widehat{g})-V(g_{\ast})|$ is bounded by $|U_{i}|$. The third inequality holds by Lemma A.1, the fourth inequality holds by the margin condition in Assumption 6.2, and the last equality holds by Assumption 6.1 or Theorem 5.1 (if $\widehat{g}$ is our surrogate optimal policy estimator). Similarly, using $g_{k}$ in the place of $\widehat{g}$, we can show that $|V(g_{k})-V(g_{\ast})|=o_{p}(n^{-\frac{1}{2}})$ because $\sup_{x}|g_{k}(x)-g_{\ast}(x)|\leq B_{k}=o(n^{-1/2})$ by Assumption 6.1. Therefore, we can conclude $|V(\widehat{g})-V(g_{k})|\leq|V(\widehat{g})-V(g_{\ast})|+|V(g_{\ast})-V(g_{k})|=o_{p}(n^{-1/2})$.
@@ A.4 Proofs for Section 6: Estimation and Inference for the Optimal Welfare
Since the identified surrogate policy function $g_{\ast}$ belongs to the set of optimal policy functions by Theorem 4.1, the identification error (A.32) of the surrogate loss approach is exactly zero. (A.29) and (A.31) are negligible by the analysis on term A in Proposition 2, respectively. It remains to further analyze the second right-hand side term (A.30), which can be decomposed as
@@
Proof of Proposition 1: The condition $k^{4+\varepsilon}/n\lesssim 1$ implies that the rate restriction $\xi_{k}\sqrt{\log(k)/n}\lesssim 1/n^{1/4}$. Also, Assumtion 5.1(iv) implies $B_{k}\lesssim(\log(k))^{-1}n^{-1/2}=o(n^{-1/2})$. The proof is complete. $\blacksquare$

### v2

@@ 7 Simulation Studies
The representative classification policy function is estimated using a sieve approximation based on a tensor-product Legendre polynomial basis. Specifically, for $k\in\{2,3,4\}$, we employ the tensor-product basis constructed from the orthonormal Legendre polynomials $\{p_{j}(x)\}_{j=0}^{\infty}$ of $k$-dimensions defined on the support $\mathcal{X}$. The sieve basis is given by $p_{1:k}(x_{1},x_{2})=\big\{\,p_{j}(x_{1})\,p_{l}(x_{2}):j,l\in[k]\,\big\},$ and the policy function is approximated as $g_{k}(x_{1},x_{2})=\sum_{j=1}^{k}\sum_{l=1}^{k}\beta_{jl}\,p_{j}(x_{1})p_{l}(x_{2}),$ or, equivalently, $g_{k}(x_{1},x_{2})=p_{1:k}(x_{1},x_{2})^{\prime}\beta.$
@@ 7.2 Gaussian Limit and Inference for the Optimal Welfare
Each plotted density from (a) to (d) closely resembles the standard normal distribution. In particular, the result in panel (a) for the plug-in estimated optimal welfare corroborates the theoretical property of Gaussian limit behavior established in Theorem 6.2. Panel (a) is comparable to panel (b) for the oracle counterpart, and this observation is consistent with the asymptotic negligibility of the preliminary representative classification policy estimation emphasized in Section 6.2. Although we report one set of results based on the logistic loss, these findings hold robustly across different surrogate loss functions.
@@ 9 Conclusion
In this paper, we have developed a toolkit for nonparametric uniform inference in cost-sensitive binary classification, a broad framework that encompasses maximum score estimation, utility-maximizing choice prediction, policy learning, and related problems. These settings are well known for slow convergence and nonstandard limiting behavior under parametric classification rules. Furthermore, nonparametric formulations of classification policy functions introduce the additional challenge of failures of identification.
To address these difficulties, we proposed a strictly convex surrogate loss that point-identifies a representative nonparametric classification policy function. Estimating this representative classification policy function provides a unified path to inference for both the optimal classification rule and its associated policy value. In particular, this strategy delivers a Gaussian approximation for the optimal classification policy and a Gaussian limit for the optimal policy value, enabling standard inferential procedures and substantially simplifying empirical implementation despite the intrinsic difficulty of the underlying problem.
Simulation evidence corroborates the theoretical results, showing that the proposed methods perform well in finite samples. Finally, our application to the National JTPA Study demonstrates how the procedure can be used to recover an optimal treatment assignment rule and to evaluate its welfare implications relative to alternative policies at hand.
While the primary contributions of this paper lie in developing inference methods for nonparametric classification policies, our approach also provides, as a by-product, practical benefits for estimating these policies. This improvement arises from two sources. First, the use of a strictly convex surrogate loss facilitates routine numerical optimization for sieve estimation. Second, and more importantly, our point identification of the representative policy function enables a straightforward estimation of the resulting classification rules.
@@ A.1 Proof for Section 4: Surrogate Identification
$C_{1}(x)=\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}=x]\quad\text{and}\quad C_{0}(x)=\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}=x].$
@@
Note that $g_{\lambda}\in\mathcal{G}$ because $\mathcal{G}$ is convex. Since $\phi$ is strictly convex, $\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}]>0$ a.s., and $\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}]>0$ a.s.,
$\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}]\cdot\phi(-g_{\lambda}(X_{i}))+\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}]\cdot\phi(g_{\lambda}(X_{i}))$
$\leq$ $\lambda[\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}]\cdot\phi(-g_{1}(X_{i}))+\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}]\cdot\phi(g_{1}(X_{i}))]$
$+(1-\lambda)[\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}]\cdot\phi(-g_{2}(X_{i}))+\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}]\cdot\phi(g_{2}(X_{i}))],$
@@ A.3 An Auxiliary Result Related to Section 6
$C_{1}(x)=\mathbb{E}[\psi^{+}(Z_{i},\eta_{0})|X_{i}=x]\quad\text{and}\quad C_{0}(x)=\mathbb{E}[\psi^{-}(Z_{i},\eta_{0})|X_{i}=x].$
@@
Suppose Assumptions 5.3, 6.1, 6.2.(ii), and conditions of Theorem 4.1(ii) hold. Then, there exists small $\delta$ such that for any $\varepsilon<\delta$, $\mathbf{1}\{|g_{\ast}(x)|\leq\varepsilon\}\leq\mathbf{1}\{|C_{1}(x)-C_{0}(x)|\leq K_{1}\varepsilon\}$ for a positive $K_{1}$ that does not depends on $x$ and $\varepsilon$.
@@
In addition to the conditions for Theorem 5.1, suppose that that Assumptions 6.1 and 6.2 hold and the sieve order for the estimator $\widehat{g}$ satisfies $k^{4+\varepsilon}/n\lesssim 1$ for some $\varepsilon>0$. Then, we have
@@
We now analyze the second term on the right-hand side. Note $|V(\widehat{g})-V(g_{k})|\leq|V(\widehat{g})-V(g_{\ast})|+|V(g_{\ast})-V(g_{k})|$. Let $U_{i}:=\mathbb{E}[\psi_{1}(Z_{i},\eta_{0})|X_{i}]-\mathbb{E}[\psi_{0}(Z_{i},\eta_{0})|X_{i}]$. Then, with some positive constant $K_{1}$,
@@
where the first inequality holds because $\widehat{g}(X_{i})$ and $g_{\ast}(X_{i})$ disagree implies that $|g_{\ast}(X_{i})|\leq|\widehat{g}(X_{i})-g_{*}(X_{i})|$, and when they disagree, the absolute difference $|V(\widehat{g})-V(g_{\ast})|$ is bounded by $|U_{i}|$. The third inequality holds by Lemma A.1, and the fourth inequality holds by the margin condition in Assumption 6.1. The last equality holds by Theorem 5.1 because the condition $k^{4+\varepsilon}/n\lesssim 1$ implies that the rate restriction $\xi_{k}\sqrt{\log(k)/n}\lesssim 1/n^{1/4}$. Also, Assumtion 5.1(iv) implies $B_{k}\lesssim(\log(k))^{-1}n^{-1/2}=o(n^{-1/2})$. Similarly, using $g_{k}$ in the place of $\widehat{g}$, we can show that $|V(g_{k})-V(g_{\ast})|=o_{p}(n^{-\frac{1}{2}})$ because $\sup_{x}|g_{k}(x)-g_{\ast}(x)|\leq B_{k}=o(n^{-1/2})$ again by Assumption 5.1(iv). Therefore, we can conclude $|V(\widehat{g})-V(g_{k})|\leq|V(\widehat{g})-V(g_{\ast})|+|V(g_{\ast})-V(g_{k})|=o_{p}(n^{-1/2})$.
@@ A.4 Proofs for Section 6: Estimation and Inference for the Optimal Welfare
Since the identified representative classification policy function $g_{\ast}$ belongs to the set of optimal policy functions by Theorem 4.1, the identification error (A.32) of the surrogate loss approach is exactly zero. (A.29) and (A.31) are negligible by the analysis on term A in Proposition 1, respectively. It remains to further analyze the second right-hand side term (A.30), which can be decomposed as